OpenStudy (anonymous):

find f(x) and g(x) such that h(x)= (f^o g)(x) h(x)=/8x+12/

OpenStudy (zzr0ck3r):

\(h(x) = 8x+12\)?

OpenStudy (anonymous):

h(x)=/8x+12/

OpenStudy (anonymous):

absolute value of 8x+12

OpenStudy (zzr0ck3r):

oh ok so if we let \(f(x) = |x|\) then we want \((f\cdot g)(x)=f(g(x)) = |g(x)| = |8x+12|\) So what is \(g(x)\) ?

OpenStudy (zzr0ck3r):

\(\huge |g(x)| = |8x+12|\) What does \(g(x)\) need to be to make this true @adaobi

OpenStudy (anonymous):

h(x)?

OpenStudy (zzr0ck3r):

I will give you this one so that you see how its done, as its hard to explain without showing you... \(g(x) = 8x+12\)

OpenStudy (anonymous):

/8x+12/= x?

OpenStudy (zzr0ck3r):

now note: \(f(x) = |x|\\\color{blue}{g(x) = 8x+12} \) so \((f\cdot g)(x)=f(\color{blue}{g(x)})=|\color{blue}{g(x)}|=|\color{blue}{8x+12}|=h(x)\)

OpenStudy (anonymous):

That is it?

OpenStudy (zzr0ck3r):

yes....

OpenStudy (anonymous):

Hmmm, thanks!